# Feature Coding in the Era of Large Models: Dataset, Test Conditions, and Benchmark

### v1

@@ 6.2 Rate-accuracy analysis
6.3 Distortion-accuracy analysis
In this subsection, we examine the linear correlation between feature MSE and task accuracy. We present the coefficients of determination in Table 5. For the VTM baseline, only depth estimation shows a high linear correlation. We believe this arises from its accuracy is measured by RMSE. However, under the Hyperprior baseline, depth estimation does not exhibit a high correlation because of a significant accuracy drop at the lowest bitrate. The common sense reasoning task exhibits an extremely low linear correlation at both baselines, indicating the different characteristics of textual features. This analysis further supports the conclusion that feature MSE is not a good metric for measuring semantic distortion. Developing effective metrics for semantic distortion remains a critical topic in feature coding.
@@ 6.3 Distortion-accuracy analysis
6.4 Ablation on truncation and quantization
In this subsection, we examine the impact of truncation and quantization on feature coding. Given the VTM baseline’s stability and deterministic properties, we conduct this ablation study on it specifically. As shown in Table 6, truncation and quantization individually result in little task accuracy loss. When both are applied together, task accuracies experience slight decreases. However, applying only quantization leads to a significant accuracy drop.
Without truncation, feature values are often limited to a small subset within the full range of [0, 1023]. In this scenario, during feature coding, the codec’s internal quantization maps the narrow input range to a single value, introducing significant distortion. However, by applying truncation, this range is expanded, helping to reduce distortion caused by the inner quantization. To demonstrate this effect, we test several truncation ranges and encode the truncated features using fixed QPs (27, 32, 37). As presented in Table 7, smaller truncation ranges generally yield higher accuracy but come at the cost of increased bitrates. Therefore, selecting an appropriate truncation range is crucial for balancing the target bitrates and accuracy.
7 Conclusion
In this paper, we highlight the critical role of feature coding in deploying large models and make three contributions to advance this field. First, we construct a comprehensive dataset that covers a wide variety of feature types. Second, we establish unified test conditions to support standardized evaluations and ensure fair comparisons for future studies. Third, we introduce two baseline methods, providing foundational benchmarks and insights for researchers exploring this field.
We identify three promising directions for future feature coding research. First, further improvements in coding performance are encouraged. Second, robust methods for measuring semantic distortion in feature coding deserve more attention. Third, developing generic feature coding methods that can handle diverse features across various models is expected.
Supplementary Material
8 Additional dataset analysis
@@
Proposed Features Existing Features
Arc-Challenge $SP_{GS}$ 27.50 -26.38 0.0038 419.14 33.28 / / / / / /
Captions from COCO2017 $SP_{H}$ 2.19 -1.09 0.5420 22918.26 118.22 / / / / / /
We visualize the distributions, feature blocks, and DCT blocks of $F_{Cls}$ and $F_{Seg}$ extracted from $SP_{DS}$ in Fig. 6 and Fig. 7. These distributions exhibit high similarities to the features extracted from $SP_{DM4}$, as they are derived from the same layer. To avoid redundancy in the main paper, these visualizations have been moved to the supplementary material.
Additionally, we compare the proposed features with existing features in terms of statistical characteristics, as presented in Table 8. Specifically, we randomly sample 10 examples from each dataset for statistical analysis, reporting the maximum, minimum, and mean values. In general, the proposed features exhibit broader distribution ranges compared to existing features. For multi-scale features, the distribution range expands with deeper network layers, whereas existing features show smaller variation. In addition, although the proposed features are more asymmetric, their mean values are closer to zero.
To assess spatial redundancy, we compute intensity variance (IV) and gradient magnitude (GM), with gradients calculated using the Sobel operator. Prior to computation, pixel and feature values are quantized to 10-bit integers. The intensity variance of the proposed features spans a wide range, from 9.61 to 22918.26, whereas existing features have a narrower range of [7323.72, 9839.49]. A similar trend is observed in gradient magnitude. The larger variations demonstrate the higher diversity of the proposed features. Furthermore, among the proposed features, textual features reveal distinct distribution characteristics compared to visual features. For instance, $SP_{GS}$ and $SP_{H}$ exhibit larger intensity variance and gradient magnitude. This highlights the necessity and importance of including textual features in the proposed dataset.
@@
Dpt NYUv2 20000 $256\times 256$ 200 1e-4 0.001, 0.005, 0.02, 0.05, 0.12
CSR Arc-Challenge, OpenBookQA 6000 $64\times 4096$ 200 1e-4 0.01405, 0.0142, 0.015, 0.045, 10
Table 9: Training details of the Hyperprior baseline.
Image Classification
Semantic Segmentation
Depth Estimation
Common Sense Reasoning
Text-to-image Snythesis
$\lambda$-BPFP
Figure 8: Visualization of rate-accuracy performance and $\lambda$-BPFP curves.
9 Baseline methods
9.1 Packing details
9.1.1 Semantic segmentation
In semantic segmentation, an image is pre-processed into multiple $518\times 518$ patches by crop and resize operations. These patches are fed into DINOv2, which produces features of shape $1370\times 1536$. In the proposed dataset, each image is pre-processed into 2 patches, resulting in features ($F_{Seg}$) with a shape of $2\times 1370\times 1536$. We reshape $F_{Seg}$ from $2\times 1370\times 1536$ into $2740\times 1536$ by vertically stacking them.
9.1.2 Depth estimation
In depth estimation, an image is first horizontally flipped. Both the original and flipped images are padded to dimensions that are multiples of 14 before being processed by DINOv2. For each image, DINOv2 generates $4\times 1611\times 1536$ embedding features at the split point $SP_{DM}$. Thus, the initial shape of $F_{Dpt}$ is $2\times 4\times 1611\times 1536$, where the two dimensions correspond to the original and flipped images. Packing $F_{Dpt}$ is performed in two steps. First, each $4\times 1611\times 1536$ feature is horizontally packed into a $1611\times 6144$ feature by concatenating the features along the width. Second, the two $1611\times 6144$ features (from the original and flipped images) are vertically stacked, resulting in a final feature shape of $3222\times 6144$.
9.1.3 Text-to-image synthesis
9.2 Training details
9.2.1 Feature Pre-processing
Interestingly, we encountered the storage problem described in the main text during the training of feature coding models. For example, just 20000 original $F_{Dpt}$ features occupy 1474 GB of storage. As the need for additional training data grows, storage demands could escalate significantly. This often exceeds the resources available to researchers in academic institutions.
To address this issue, we crop the features into smaller dimensions to facilitate feature coding training. The specific cropping configurations are detailed in Table 9. However, it is important to note that vertical cropping is not effective for $F_{CSR}$. In our experiments, when $F_{CSR}$ is cropped into $64\times 512$, training becomes unstable, and the coding performance deteriorates significantly. We hypothesize that this limitation arises because $F_{CSR}$ is extracted from an LLM and involves only textual data.
It is worth emphasizing that while cropping is effective for feature coding training, this strategy is not applicable to large model training. For large models, it is unreasonable to crop portions of features to accelerate training, as downstream models typically require the entire semantic information. For example, cropping a $256\times 256$ patch from $F_{Dpt}$ would make it challenging to perform accurate depth estimation. Therefore, designing efficient feature coding methods remains crucial for reducing the storage burden in large model deployments.
Additionally, truncation and normalization operations are applied. The truncation regions listed in Table 4 are used. After truncation, all features are normalized into [0, 1].
9.2.2 Loss Function
In our experiments, we train models separately for each task. Across all tasks, the same loss function is employed. The loss function comprises two components: bitrate and feature MSE, defined as:
$L=BPFP+\lambda\times||(F_{o}-\hat{F_{r}})||^{2}$ (1)
where $\lambda$ is a scaling factor used to adjust the bitrate. A larger $\lambda$ corresponds to a higher bitrate.
9.2.3 Hyperparameter Setting
The number of training epochs, initial learning rate, and $\lambda$ values are provided in Table 9. The number of training epochs is determined based on the size of the training dataset. In our experiments, all training processes converge to a stable loss. Five different $\lambda$ values are used to achieve various bitrates. The $\lambda$ values are selected to align the bitrates with those achieved by the VTM baseline.
Image Classification
Image Classification
Semantic Segmentation
Semantic Segmentation
Depth Estimation
Depth Estimation
Common Sense Reasoning
Common Sense Reasoning
Text-to-Image Synthesis
Text-to-Image Synthesis
Figure 9: Visualization of the linear relationship between feature MSE and task accuracy. Left: the VTM baseline. Right: the Hyperprior baseline.
Image Classification Semantic Segmentation
BPFE Accu. Accu. Drop BPFE mIoU mIoU Drop
VTM Baseline 1.90 100 0.00% 1.78 79.68 2.35%
0.98 99 1.00% 0.90 78.91 3.30%
0.21 81 19.00% 0.24 73.53 9.89%
0.04 18 82.00% 0.04 55.05 32.54%
0.01 2 98.00% 0.01 32.64 60.00%
Hyperprior Baseline (Trained on Classification) 1.99 92 8.00% 1.53 77.75 4.72%
1.11 91 9.00% 0.88 75.22 7.82%
0.89 86 14.00% 0.70 72.88 10.68%
0.37 29 71.00% 0.24 62.23 23.73%
0.23 11 89.00% 0.09 48.24 40.89%
Hyperprior Baseline (Trained on Segmentation) 2.41 97 3.00% 1.72 78.44 3.87%
1.74 97 3.00% 1.30 77.85 4.59%
0.87 86 14.00% 0.54 76.39 6.39%
0.25 20 80.00% 0.12 62.69 23.17%
0.09 1 99.00% 0.03 40.92 49.85%
Table 10: Rate-accuracy performance comparisons among different feature coding models.
10 Additional experimental results
10.1 Rate-accuracy analysis
We present the rate-accuracy curves in Fig. 8 to give a more straightforward comparison across different tasks. For semantic segmentation and depth estimation, the Hyperprior baseline achieves comparable coding performance to the VTM baseline. However, for the other three tasks, the performance is noticeably poorer. In the case of text-to-image synthesis, this is expected because $F_{TTI}$ shares a distribution similar to that of images and VTM has a higher coding performance in images than the Hyperprior codec. For image classification and common sense reasoning, improved truncation regions or quantization methods could potentially enhance coding performance. However, as feature pre-processing is not the focus of this paper, we leave this exploration to future studies.
For the VTM baseline, we conducted an ablation study on truncation regions. VTM-base, VTM-trun2, and VTM-trun3 correspond to the truncation regions specified in Table 4, the left one of Table 7, and the right one of Table 7, respectively. For depth estimation and common sense reasoning, the tested truncation regions significantly impact coding performance. As with feature pre-processing, the exploration of optimal truncation regions is deferred to future work.
10.2 Distortion-accuracy analysis
We visualize the distortion-accuracy curves in Fig. 9. The VTM baseline performs best in depth estimation, while the Hyperprior baseline achieves the highest linear correlation in text-to-image synthesis. These different behaviors highlight the distinct coding characteristics of the two methods. However, both baselines demonstrate the ineffectiveness of feature MSE as a metric for measuring semantic distortion.
$SP_{DM1},QP=42$
$SP_{DM1},\lambda=0.001$
$SP_{DM2},QP=42$
$SP_{DM2},\lambda=0.001$
$SP_{DM3},QP=42$
$SP_{DM3},\lambda=0.001$
$SP_{DM4},QP=42$
$SP_{DM4},\lambda=0.001$
$SP_{GS},QP=32$
$SP_{GS},\lambda=0.01405$
$SP_{H},QP=42$
$SP_{H},\lambda=0.005$
10.3 Exploration of generalization ability
We verify the generalization ability of the two baseline methods. First, we examine the $\lambda$-BPFP relationship and present the results in the bottom-right corner of Fig. 8. The analysis reveals that the same $\lambda$ produces different rate-distortion trade-offs across various tasks, indicating that features extracted from different tasks do not share similar coding properties.
To further investigate this, we select the two most similar tasks (image classification and semantic segmentation) based on their $\lambda$-BPFP curves and compare their rate-accuracy performance, as shown in Table 10. For the VTM baseline, although similar bitrates are achieved, the accuracy drop varies significantly at low bitrates. For the Hyperprior baseline, we encode $F_{Cls}$ and $F_{Seg}$ using two groups of models: models trained for image classification and models trained for semantic segmentation. Regardless of which group of models is used, different accuracy drops are observed.
These analyses indicate that the two proposed baselines lack generalization ability across various tasks. The generalization ability of feature coding methods across multiple tasks has received little attention in prior research. By presenting these results, we aim to draw more attention to this important research direction.
10.4 Feature distribution analysis
We visualize the distribution of reconstructed features in Fig. 10, comparing the two baseline methods at similar bitrates. For $SP_{DM}$, the Hyperprior baseline fails to reconstruct features within their original distribution range. The reconstructed features appear truncated during the feature coding process. We attribute this truncation to the inner quantization (rounding operation) in the Hyperprior codec, where the transformed latent variables are directly rounded to integers before entropy coding. In contrast, the VTM baseline quantizes original features into a broader distribution range of [0, 1023], effectively preventing the truncation phenomenon. A similar truncation phenomenon is observed for $SP_{H}$. For $SP_{GS}$, we visualize features encoded at a higher bitrate and observe that the truncation phenomenon disappears.
This truncation phenomenon highlights the incompatibility between features from large models and learned image coding methods. Two promising research directions emerge: (1) designing a new framework specifically tailored for feature coding, and (2) developing pre-processing techniques to adapt original features for compatibility with image coding methods.

### v3

@@ 9.3 Ablation on truncation and quantization
9.4 Distribution analysis on decoded features
We visualize the distribution of reconstructed features in Fig. 7, comparing the two baseline methods at similar bitrates. For $SP_{DM}$, the Hyperprior baseline fails to reconstruct features within their original distribution range. The reconstructed features appear truncated during the feature coding process. We attribute this truncation to the Hyperprior’s internal quantization (rounding operation), where the transformed latent variables are directly rounded to integers before entropy coding. In contrast, the VTM baseline quantizes original features into a broader distribution range of [0, 1023], effectively preventing the truncation phenomenon. A similar truncation phenomenon is observed for $SP_{H}$. For $SP_{GS}$, the features are encoded at a higher bitrate and the truncation phenomenon is not observed.
10 Additional dataset analysis
In addition to the visualizations provided in the main text, we present further statistical analyses of the proposed dataset.
10.1 Distribution analysis
We compare the statistical characteristics of the proposed features with those of existing features by randomly sampling 10 examples from each dataset and reporting their maximum, minimum, and mean values in Table 9. Overall, the proposed features exhibit a broader distribution range compared to existing features. In particular, for multi-scale features, the distribution range expands in deeper network layers, whereas existing features exhibit smaller variations. Moreover, although the proposed features tend to be more asymmetric, their mean values remain closer to zero.
@@ 10.1 Distribution analysis
10.2 Redundancy analysis
To evaluate spatial redundancy, we compute intensity variance (IV) and gradient magnitude (GM), where gradients are derived using the Sobel operator. Before these computations, both pixel and feature values are quantized to 10-bit integers. The intensity variance of the proposed features spans a wide range, from 9.61 to 22918.26, whereas existing features have a narrower range from 7323.72 to 9839.49. A similar trend is observed in gradient magnitude. These broader variations underscore the higher diversity of the proposed features. Furthermore, among the proposed features, textual features reveal distinct distribution characteristics compared to visual features. For instance, $SP_{GS}$ and $SP_{H}$ exhibit larger intensity variance and gradient magnitude. This observation highlights the necessity and importance of incorporating textual features in the dataset.
10.3 Visualization
We visualize the distributions, feature blocks, and DCT blocks of $F_{Cls}$ and $F_{Seg}$ extracted from $SP_{DS}$ in Fig. 8 and Fig. 9. These distributions exhibit high similarities to the features extracted from $SP_{DM4}$, as they are derived from the same layer. To avoid redundancy in the main paper, these visualizations have been moved here.